# DeepSport: A Multimodal Large Language Model for Comprehensive Sports Video Reasoning via Agentic Reinforcement Learning

### v1

@@ 3.3 Two-Stage Training Strategy
We first perform Supervised Fine-Tuning (SFT) to teach the model the CoT format as the “cold start”, which has been wildly adopted in recent works [deepseekai2025deepseekr1incentivizingreasoningcapability, feng2025video, zhang2025thinkingvideosmultimodaltoolaugmented], followed by a Reinforcement Learning (RL) phase to optimize the model’s reasoning and tool-use policy.
3.3.1 SFT Cold-Start
We first perform an SFT phase using the DeepSport-CoT-15k. The objective of this phase is not to teach the model to solve all tasks, but to serve as a “cold-start” to familiarize the model with the complex CoT trajectory structure. By training on the DeepSport-CoT-15k, the model learns the syntax of our thought and action tags (e.g., <think>... and <tool_call>...) and the basic patterns of multi-turn reasoning. This initial alignment is crucial for stabilizing the subsequent RL phase, as it ensures the model starts with a reasonable policy capable of producing valid, well-formatted actions to explore.
@@ 3.3.2 Agentic RL with Gated Tool Rewards
During the reinforcement learning phase, we employ the Group Relative Policy Optimization (GRPO) [shao2024deepseekmath], which has been commonly used in reinforcement learning for model training across multiple domains [deepseekai2025deepseekr1incentivizingreasoningcapability, liu2025finr1largelanguagemodel, feng2025video]. For each prompt $q$ from the RL set, we sample a group of $G$ trajectories $\{\tau_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}(\cdot\mid q)$ and obtain rewards $\{r_{i}\}_{i=1}^{G}$. The advantage $A_{i}$ is defined by the group statistics:
@@
To train our model, we design a novel, multi-component reward function $R(\tau)$ that provides granular behavioral shaping to encourage both task success and effective reasoning.
@@
1. Core Objective. The primary reward signal $R_{\text{acc}}$ measures task success, based on the semantic accuracy of the model’s final answer in trajectory $\tau$:
$R_{\text{acc}}(\tau)=acc(\tau),\quad acc(\tau)\in[0,1].$ (4)
@@
Table 2: Performance comparison on the test set. We evaluate models across four high-level dimensions designed to assess holistic sports tasks. Each dimension encompasses diverse specific capabilities: Fine-Grained Recognition involves visual perception tasks (e.g., Action Recognition, Player Identification); Rule & Procedural tests regulatory knowledge (e.g., Foul Classification, Referee Decision Analysis); Assessment & Coaching evaluates expert reasoning (e.g., Tactical Analysis, Scoring Prediction); and Live Commentary & Reporting assesses multimodal narration (e.g., Play-by-Play Commentary, Score Reading).
@@
Qwen2.5-VL-7B-Instruct 16 28.20 13.71 9.09 3.13 16.98
DeepSport (Ours) 14.39 51.09 43.82 24.74 24.60 40.08
2. Behavioral Shaping. We introduce shaping rewards controlled by gating variables. We define the tool usage gate $g_{\text{tool}}(\tau)\in\{0,1\}$, which activates ($g_{\text{tool}}=1$) if the model invokes the frame_extraction_tool at least once. The accuracy gate $g_{\text{acc}}(\tau)$ activates if the accuracy is above a minimum threshold (e.g., $acc(\tau)\geq 0.5$):
$g_{\text{acc}}(\tau)=\begin{cases}1,&\text{if }acc(\tau)\geq 0.5,\\ 0,&\text{otherwise.}\end{cases}$ (5)
These gates define the tool usage incentive $R_{\text{tool}}$, which is designed to (a) strongly reward successful tool use, (b) provide a small “curiosity” bonus for exploratory tool use, and (c) give no reward for failing to use tools:
$R_{\text{tool}}(\tau)=\begin{cases}0.5\cdot acc(\tau),&\text{if }g_{\text{tool}}(\tau)=1\text{ and }g_{\text{acc}}(\tau)=1,\\[4.0pt] 0.03,&\text{if }g_{\text{tool}}(\tau)=1\text{ and }g_{\text{acc}}(\tau)=0,\\[4.0pt] 0,&\text{if }g_{\text{tool}}(\tau)=0.\end{cases}$ (6)
3. Format Gating and Final Reward. Finally, we enforce the structural and logical integrity of the trajectory (defined in Section 3.1). We use a format gate $g_{\text{fmt}}(\tau)\in\{0,1\}$ that validates the <tool_call>...</tool_call> syntax and penalizes invalid actions, such as redundant tool calls on the exact same interval. Invalid trajectories incur a structural penalty:
$P_{\text{format}}(\tau)=-0.05\,(1-g_{\text{fmt}}(\tau)).$ (7)
The total reward is then expressed as:
$R(\tau)=g_{\text{fmt}}(\tau)\big(R_{\text{acc}}(\tau)+R_{\text{tool}}(\tau)\big)+P_{\text{format}}(\tau).$ (8)
When the output format is valid ($g_{\text{fmt}}=1$), the reward combines task accuracy and tool-use behavior. Otherwise ($g_{\text{fmt}}=0$), the reward collapses to the penalty $R(\tau)=-0.05$.
4 Experiments
4.1 Experimental Setup
Datasets.
As described in Section 3.2, our data pipeline yields three distinct subsets based on a strict video-level split. First, we construct DeepSport-CoT-15k, consisting of 15k high-quality, LLM-filtered CoT QA pairs used specifically for the SFT cold-start phase. Second, for the reinforcement learning phase, we utilize DeepSport-RL-63k, which comprises 63k original, unified QA pairs (prompt plus ground truth answer) serving as the environment prompts. Finally, we evaluate performance on our held-out Testing Benchmark, containing 6.7k QA pairs that cover all tasks and sports to ensure a rigorous evaluation.
Implementation Details.
We select Qwen2.5-VL-7B [bai2025qwen2] as our backbone model, processing video inputs at a resolution of $640\times 360$. Our training pipeline consists of two stages: we first fine-tune the model on the DeepSport-CoT-15k set for 1 epoch, followed by reinforcement learning via our tool-based GRPO algorithm with a batch size of 32 and a rollout number of 8 on the DeepSport-RL-63k set for 300 steps. The model is trained to generate up to 12,800 tokens per response. Training was performed on 8 $\times$ H20 GPUs for totaling 796 GPU hours.
Baselines.
We compare DeepSport against a comprehensive suite of MLLMs, including both open- and closed-source models. Specifically, we select models from the Qwen family (Qwen2.5-VL-7B-Instruct as our primary baseline for direct comparison, Qwen3-VL-8B-Instruct/Thinking, and Qwen3-VL-235B-A22B-Thinking [bai2025qwen2, qwen3technicalreport]), as well as InternVL3.5-14B [wang2025internvl3_5] and Video-R1 [feng2025video]. For evaluation, we uniformly sample 16 frames per video.
Figure 3: We present a comparison on the diving code detection task, where the model need capture each fine-grained movement. The Qwen2.5-VL-7B-Instruct model, relying on passive, single-pass processing of 16 sparsely sampled frames, misses the high-speed contact and made the wrong conclusion. Our DeepSport model, despite having only 7B parameters, active the multi-turn conversation by involking the frame_extraction_tool(35, 59) to retrieve second roun relevant frames. Using this new evidence, it correctly identifies the diving code successfully. This demonstrates the superiority of our active, iterative reasoning paradigm over static models.
4.2 Main Results
We evaluate all models on our 6.7k multi-task, multi-sport test set using an LLM-as-Judge (Claude 4.5 Sonnet [anthropic_2025_introducing]), with scores scaled to a percentage format (0-100).
The comprehensive evaluation results are presented in Table 2. This trend of high performance extends across the diverse task spectrum of our benchmark. When viewed holistically, DeepSport’s well-rounded capabilities become clear. It achieves an overall average score of 40.08, establishing a new state-of-the-art by surpassing the powerful closed-source GPT-5 and the massive Qwen3-VL-235B-A22B-Thinking.
@@ 4.2 Main Results
Quantitatively, DeepSport’s superiority is most pronounced in perception-heavy and rule-bound tasks. It secures a dominant lead in Fine-Grained Recognition with a score of 51.09 and Rule & Procedural tasks with 43.82. This substantial margin over generalist models validates our core hypothesis: the active tool-use paradigm effectively captures fleeting visual details missed by passive processors, while agentic RL successfully internalizes complex domain-specific regulations.
@@
In terms of efficiency, while the Qwen3-VL-235B-A22B-Thinking model retains a marginal edge in generative tasks like Assessment & Coaching and Live Commentary—likely leveraging its massive pre-trained knowledge base—DeepSport delivers highly competitive performance despite having only $\sim$3% of the parameter count. This signifies that our distillation pipeline effectively transfers high-level reasoning capabilities to a compact backbone. Moreover, our model achieves the SOTA overall score using only 14.39 frames on average, fewer than the fixed 16 frames used by other models.
@@
Furthermore, the necessity of specialized training is underscored by the comparison with general-domain baselines. DeepSport achieves a score improvement of 23.1 over its own backbone, Qwen2.5-VL-7B, and significantly outperforms the general RL-trained Video-R1 model. This performance gap underscores that general-domain training fails to capture sport-specific nuances, highlighting the critical need for our multi-task specialized model. Figure 3 shows the output comparison between our DeepSport model and the backbone model, Qwen2.5-VL-7B.
@@ 4.3 Generalization on General Benchmarks
To investigate whether our sport-specific training compromises general video understanding capabilities, we evaluated DeepSport on two widely used general benchmarks: MLVU [zhou2024mlvu] and Video-MME (Long) [fu2024video]. As shown in Table 3, DeepSport exhibits a slight performance drop compared to its backbone, Qwen2.5-VL-7B-Instruct. This marginal degradation is a common trade-off in domain-specific fine-tuning, which has been discussed in luo2025empirical work. However, notably, DeepSport achieves this competitive performance using significantly fewer frames ($\sim$18 frames) compared to the fixed 32 frames of the backbone. This demonstrates that our active frame extraction mechanism generalizes beyond sports, allowing the model to efficiently locate key information in general long videos while maintaining robust performance superior to previous baselines like VideoChat2 [li2023videochat].
Table 3: Performance on general video understanding benchmarks. DeepSport maintains competitive results with lower average frame usage.
Model MLVU Test Video-MME (Long)
Score Avg. Frames Score Avg. Frames
VideoChat2 35.10 16.00 33.20 16.00
Qwen2.5-VL-7B-Instruct 41.60 32.00 41.90 32.00
DeepSport (Ours) 39.62 18.97 39.98 18.37
@@ 4.4 Error Analysis
Figure 4: Error analysis on 70 sampled failure cases. The dominance of Tool Grounding Failure (42.9%) highlights that precise temporal localization remains the primary bottleneck, followed by fine-grained Visual Hallucination (37.1%).
As shown in Figure 4, the dominance of Tool Failure indicates that while our model is capable of reasoning, the primary bottleneck remains accurate temporal localization in long videos. Furthermore, the high rate of Visual Hallucination suggests that future work should focus on enhancing the fine-grained visual encoder resolution.

### v2

@@ 3.3 Two-Stage Training Strategy
We first perform Supervised Fine-Tuning (SFT) to teach the model the CoT format as the “cold start”, which has been wildly adopted in recent works [10, 14, 72], followed by a Reinforcement Learning (RL) phase to optimize the model’s reasoning and tool-use policy.
3.3.1 Sports Curriculum SFT Cold-Start
We first perform a Supervised Fine-Tuning (SFT) phase using the DeepSport-CoT-14K dataset. The objective of this phase is twofold: (1) to serve as a “cold-start” to familiarize the model with the complex CoT trajectory structure (i.e., learning the syntax of <think>... and <tool_call>...), and (2) to incrementally build the model’s domain knowledge.
To achieve this, we introduce a Sports Curriculum Learning strategy. The underlying rationale is that high-level analytical tasks (e.g., Live Commentary, Assessment & Coaching) are fundamentally predicated on solid foundational perception (Fine-Grained Recognition). Therefore, we partition the SFT samples into 5 progressive stages. As training advances from Stage 1 to Stage 5, we gradually decrease the proportion of Fine-Grained Recognition tasks (from 79% down to 47%) while steadily increasing the ratio of complex reasoning tasks, such as Rule & Procedural Logic (from 7% to 20%) and Assessment & Coaching (from 5% to 21%). This progressive data curriculum ensures that the model masters basic visual grounding before tackling expert-level logic. This initial alignment is crucial for stabilizing the subsequent RL phase, providing a reasonable policy capable of producing valid actions.
@@ 3.3.2 Agentic RL with Gated Tool Rewards
During the reinforcement learning phase, we employ the Group Relative Policy Optimization (GRPO) [47], which has been commonly used in reinforcement learning for model training across multiple domains [10, 32, 14]. For each prompt $q$ from the RL set, we sample a group of $G$ trajectories $\{\tau_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}(\cdot\mid q)$ and obtain rewards $\{r_{i}\}_{i=1}^{G}$. The advantage $A_{i}$ is defined by the group statistics:
@@
To train our model, we design a multi-component reward function $R(\tau)$ that encourages task success, selective tool invocation, and well-formed outputs.
1. Sample Classification. A naïve tool-use bonus would incentivize the model to invoke the frame_extraction_tool indiscriminately, even when the initial frames already suffice. To prevent this reward hacking, we pre-classify every training sample by evaluating the base model on the initial uniformly-sampled frames before RL training:
  • Class-A: samples the base model fails to answer correctly from the initial frames, indicating that temporal re-inspection is likely necessary.
  • Class-B: samples the base model can answer correctly from the initial frames, indicating that further tool invocation is unnecessary.
We denote the class of a trajectory $\tau$ as $c(\tau)\in\{\texttt{a},\texttt{b}\}$.
@@
2. Core Objective. The primary reward signal $R_{\text{acc}}$ measures task success via an LLM-as-a-judge semantic similarity score:
$R_{\text{acc}}(\tau)=\mathrm{acc}(\tau),\quad\mathrm{acc}(\tau)\in[0,1].$ (4)
3. Selective Tool-Use Incentive. For Class-A samples (initially unanswerable), the model is rewarded for using the tool when it leads to a correct answer; for Class-B samples (initially answerable), the model is rewarded for refraining from tool use while still answering correctly:
$R_{\text{tool}}(\tau)=\begin{cases}\dfrac{\mathrm{acc}(\tau)}{2},&\text{if }c(\tau){=}\texttt{a},\;\text{used tool},\;\mathrm{acc}(\tau){\geq}0.5,\\ \dfrac{\mathrm{acc}(\tau)}{2},&\text{if }c(\tau){=}\texttt{b},\;\text{no tool},\;\mathrm{acc}(\tau){\geq}0.5,\\ 0,&\text{otherwise.}\end{cases}$ (5)
This asymmetric design closes the reward hacking loophole: the model cannot earn $R_{\text{tool}}$ by blindly invoking the tool on every sample, as doing so on Class-B samples yields zero bonus. Instead, the model must learn to judge whether the initial frames are sufficient before deciding to act.
4. Format Gating and Total Reward. We enforce the structural integrity of the trajectory via a format validation that checks matched and correctly nested <think>, <tool_call>, and <answer> tags, non-empty tag contents, and valid tool-call arguments (e.g., strictly narrowing frame ranges with no redundant re-requests). The total reward is:
$R(\tau)=\begin{cases}R_{\text{acc}}(\tau)+R_{\text{tool}}(\tau),&\text{if format is valid,}\\[4.0pt] -0.05,&\text{otherwise.}\end{cases}$ (6)
4 Experiments
4.1 Experimental Setup
Datasets.
As described in Section 3.2, our data pipeline yields three distinct subsets based on a strict video-level split. First, we construct DeepSport-CoT-14k, consisting of 14k high-quality, LLM-filtered CoT QA pairs used specifically for the SFT cold-start phase. Second, for the reinforcement learning phase, we utilize DeepSport-RL-63k, which comprises 63k original, unified QA pairs (prompt plus ground truth answer) serving as the environment prompts. Finally, we evaluate performance on our held-out Testing Benchmark, containing 6.7k QA pairs that cover all tasks and sports to ensure a rigorous evaluation.
Implementation Details.
We select Qwen2.5-VL-7B [3] as our backbone model, processing video inputs at a resolution of $640\times 360$. Our training pipeline consists of two stages: we first fine-tune the model on the DeepSport-CoT-14k set for 1 epoch following the 5-stage curriculum learning design, followed by reinforcement learning via our tool-based GRPO algorithm with a batch size of 32 and a rollout number of 8 on the DeepSport-RL-63k set for 300 steps. The model is trained to generate up to 12,800 tokens per response. Training was performed on 8 $\times$ H20 GPUs for totaling 796 GPU hours.
Baselines.
We compare DeepSport against a comprehensive suite of MLLMs using uniform 16-frame sampling per video, including both open- and closed-source models . Specifically, we select models from the Qwen family (Qwen2.5-VL-7B-Instruct as our primary baseline for direct comparison, Qwen3-VL-8B-Instruct/Thinking, and Qwen3-VL-235B-A22B-Thinking [3, 50]), as well as InternVL3.5-14B [57] and Video-R1 [14]. In addition, we evaluate DeepSport under alternative frame-sampling strategies, including T [69] and VideoITG [56], and we further compare against the soccer-specific model SoccerChat [17].
4.2 Main Results
We evaluate models on our 6.7k multi-task, multi-sport test set using an LLM-as-Judge (Claude 4.5 Sonnet [2]), with scores scaled to a percentage format.
The comprehensive evaluation results are presented in Table 2. This trend of high performance extends across the diverse task spectrum of our benchmark. When viewed holistically, DeepSport’s well-rounded capabilities become clear. It achieves an overall average score of 37.67, establishing a new state-of-the-art by surpassing the powerful closed-source GPT-5[48] and the massive Qwen3-VL-235B-A22B-Thinking.
@@ 4.2 Main Results
Table 2: Performance comparison on the test set. We evaluate models across four high-level dimensions designed to assess holistic sports tasks. Each dimension encompasses diverse specific capabilities: Fine-Grained Recognition involves visual perception tasks (e.g., Action Recognition, Player Identification); Rule & Procedural tests regulatory knowledge (e.g., Foul Classification, Referee Decision Analysis); Assessment & Coaching evaluates expert reasoning (e.g., Tactical Analysis, Scoring Prediction); and Live Commentary & Reporting assesses multimodal narration (e.g., Play-by-Play Commentary, Score Reading). Bold indicates the best performance, and blue background indicates the second best. Green arrows ($\uparrow$) denote the absolute performance gains of our DeepSport over the Qwen2.5-VL-7B baseline.
@@
T 16 35.79 19.49 9.57 7.56 22.30
VideoITG 16 35.39 15.25 9.64 5.96 20.85
SoccerChat 16 31.43 29.20 8.69 10.23 23.19
Qwen2.5-VL-7B-Instruct (Baseline) 16 28.20 13.71 9.09 3.13 16.98
Baseline + SFT (Ours) 10 40.67 24.38 21.40 21.07 29.91
Baseline + Sports Curriculum SFT (Ours) 8.72 46.11 32.61 17.87 17.70 32.48
[…]
@@
Quantitatively, DeepSport’s superiority is most pronounced in perception-heavy and rule-bound tasks. It secures a dominant lead in Fine-Grained Recognition with a score of 49.89 and Rule & Procedural tasks with 41.20. This substantial margin over generalist models validates our core hypothesis: the active tool-use paradigm effectively captures fleeting visual details missed by passive processors, while agentic RL successfully internalizes complex domain-specific regulations.
To validate our progressive training paradigm, we conducted an ablation study on the training stages (shown in the lower section of Table 2). While direct SFT yields a significant performance bump, applying our staged progressive training (Sports Curriculum SFT) further elevates the overall average score from 29.91 to 32.48. Notably, this curriculum drives massive gains in foundational perception (Fine-Grained Recognition improves from 40.67 to 46.11) and regulatory logic (Rule & Procedural improves from 24.38 to 32.61). This clearly demonstrates that mastering basic visual grounding before tackling expert-level logic is a highly effective strategy for sports video understanding.
@@
In terms of efficiency, while the Qwen3-VL-235B-A22B-Thinking model retains a marginal edge in generative tasks like Assessment & Coaching and Live Commentary—likely leveraging its massive pre-trained knowledge base—DeepSport delivers highly competitive performance despite having only $\sim$3% of the parameter count. This signifies that our distillation pipeline effectively transfers high-level reasoning capabilities to a compact backbone. Moreover, our model achieves the SOTA overall score using only 9.81 frames on average, significantly fewer than the fixed 16 frames used by other models.
@@
Furthermore, the necessity of specialized training is underscored by the comparison with general-domain baselines. DeepSport achieves a score improvement of 20.69 over its own backbone, Qwen2.5-VL-7B, and significantly outperforms the general RL-trained Video-R1 model. This performance gap underscores that general-domain training fails to capture sport-specific nuances, highlighting the critical need for our multi-task specialized model. Figure 3 shows the output comparison between our DeepSport model and the backbone model, Qwen2.5-VL-7B.
Figure 3: We present a comparison on the foul detection task, where the model need capture fine-grained movement. The Qwen2.5-VL-7B-Instruct model, relying on passive, single-pass processing of 16 sparsely sampled frames, misses the high-speed contact and made the wrong conclusion. Our DeepSport model activates the multi-turn conversation by invoking the frame_extraction_tool(30, 60) to retrieve second-round relevant frames. Our model correctly identifies the foul.
@@ 4.3 Generalization on General Benchmarks
To investigate whether our sport-specific training compromises broader video understanding capabilities, we expanded our evaluation beyond sports to encompass both general long video benchmarks and fine-grained action recognition datasets. Specifically, we evaluate DeepSport on two distinct dimensions: (1) long-form video understanding using LVBench [75], LongVideoBench [59], and Video-MME (Long) [15]; and (2) fine-grained action and motion understanding using DREAM-1K [55], MotionBench [24], and ActionAtlas [45].
As shown in Table 3, DeepSport maintains competitive performance on general video tasks. While there is a slight performance drop on VideoMME (Long) compared to its backbone, Qwen2.5-VL-7B-Instruct, a common trade-off in domain-specific fine-tuning [33], DeepSport improves upon the baseline in both LVBench and LongVideoBench. Most notably, it achieves these robust results while processing significantly fewer frames ($\sim$13 frames) compared to the fixed 32 frames of the backbone, demonstrating the cross-domain efficiency of our active frame extraction mechanism.
Furthermore, DeepSport exhibits substantial improvements in general action recognition tasks. On broader motion benchmarks like DREAM-1K and MotionBench, DeepSport significantly outperforms the baseline by effectively capturing fleeting motion dynamics through its iterative, multi-turn frame sampling. To further test the generalization limits within the sports domain, we specifically evaluated DeepSport on unseen sports in the ActionAtlas benchmark. We purposefully selected sports that were entirely excluded from our training set. DeepSport still surpasses the baseline model. This strongly indicates that our model does not merely overfit to or memorize sport-specific rules; instead, it successfully internalizes the underlying interconnectedness of athletic human movements, enabling strong transferability to novel disciplines.
Table 3: Performance on general video understanding and action recognition benchmarks. DeepSport maintains competitive performance on long videos while significantly improving general and unseen sports action recognition, all with fewer average frames.
Model General Video Understanding Action & Motion Recognition
LVBench LongVideoBench VideoMME (Long) DREAM-1K MotionBench ActionAtlas
Acc Frames Acc Frames Acc Frames F1 Frames Acc Frames Acc Frames
Qwen2.5-VL-7B-Instruct 31.6 32 43.2 32 41.6 32 26.0 16.0 43.1 16.0 25.6 16.0
DeepSport (Ours) 32.0 13.2 45.9 12.7 40.4 12.5 30.5 8.0 48.5 8.3 27.2 8.5
@@ 4.4 Error Analysis
Figure 4: Error analysis. Tool Grounding Failure highlights that precise temporal localization remains the primary bottleneck, followed by fine-grained Visual Hallucination.
As shown in Figure 4, the dominance of Tool Failure indicates that while our model is capable of reasoning, the primary bottleneck remains accurate temporal localization. Furthermore, the high rate of Visual Hallucination suggests that future work should focus on enhancing the fine-grained visual ability.